Mary is 18 years older than Gabe. 15 years ago, Mary was 4 times as old as Gabe. How old is Gabe right now?
Found 2 solutions by math_helper, stanbon:
Answer by math_helper(2461)

Let M = Mary's age
Let G = Gabe's age
M = G + 18 (1)
M-15 = 4(G-15) (2)
Subs G+18 for M from (1) into (2):
(G+18) - 15 = 4(G-15)
G + 3 = 4G - 60
-3G = -63
G = -63/-3 = 63/3 = 21
M=G+18 = 21+18 = 39

Ans:
Gabe is 21, Mary is 39

Check:
(39-15) = 24
4*(21-15) = 4*6 = 24 (ok)

Answer by stanbon(75887)

Mary is 18 years older than Gabe. 15 years ago, Mary was 4 times as old as Gabe. How old is Gabe right now?
-----
Equations:
M = G + 18
M-15 = 4(G-15)
--------
Substitute for "M" and solve for "G"::
G+18-15 = 4G-60
------
3G = 63
----
Gabe is 21 yrs old
Cheers,
Stan H.
-----------